OpenStudy (anonymous):

A segment with endpoints F(4,9) and G (7,2) is divided by a point H such that FH and GH form a 1:3 ratio. find the y value for H. a)6.75 b)7.25 c)8.25 d)5.75

OpenStudy (anonymous):

help pleeeease

OpenStudy (anonymous):

@jdoe0001 hey can you please try this one? :)

OpenStudy (jdoe0001):

so the line looks like 

OpenStudy (jdoe0001):

so.. gimme one sec

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf F(4,9)\qquad G(7,2)\qquad ratio1=1\qquad ratio2=3\qquad 1:3\\ \quad \\ \quad \\ \cfrac{FH}{GH}=\cfrac{ratio1}{ratio2}\implies ratio2\cdot FH=ratio1\cdot GH\quad \textit{dividing by H}\\ \quad \\ ratio2\cdot F=ratio1\cdot G\implies 3(4,9)=1(7,2)\\ \quad \\\qquad {\color{blue}{ H=\left(\cfrac{\textit{sum of "x" values}}{ratio1+ratio2}\quad ,\quad \cfrac{\textit{sum of "y" values}}{ratio1+ratio2}\right)}}\\ \quad \\ \qquad thus\qquad \\ \quad \\ H=\left(\cfrac{(3\cdot 4)+(1\cdot 7)}{1+3}\quad ,\quad \cfrac{(3\cdot 9)+(1\cdot 2)}{1+3}\right)\)

OpenStudy (anonymous):

ok i got (4.75, 7.5) now what do i do with that?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

so yes.. .that's correct.. so recall find the \({\color{brown}{ y\ value}}\) for H.

OpenStudy (anonymous):

so which of the answer choices?

OpenStudy (jdoe0001):

well what's the y-coordinate for H that you found?

OpenStudy (anonymous):

7.5... but the answer choice has 7.25

OpenStudy (jdoe0001):

hmmm ohh I see. . misread you a bit right... recheck your calculation for "y" :)

OpenStudy (jdoe0001):

\(\bf \cfrac{(3\cdot 9)+(1\cdot 2)}{1+3}\)

OpenStudy (anonymous):

yes lol i always mistype something on my calculator lol thanks for the help

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

sooo the answer is 7.25?